what is the answer to the problem 2x-4+x=-3x+7-10
Answer by checkley79(3341)

2x-4+x=-3x+7-10
3X-4=-3X-3
3X+3X=-3+4
6X=1
X=1/6 ANS.
PROOF:
2*1/6-4+1/6=-3*1/6+7-10
2/6-4+1/6=-3/6-3
3/6-4=-3.5
-3.5=-3.5